### v1

@@ 7 Limitation and Future Work
$\mathcal{L}[v_{\infty}]-\mathcal{L}[v_{t}]$ $<(1-\epsilon)(\mathcal{L}[\bar{v}]-\mathcal{L}[v_{\infty}]),\quad\text{where}\quad\bar{v}(s):=v_{\infty}(s)+(1-\delta)(\tau-\gamma_{A})\lVert v^{*}-v_{\infty}\rVert_{\infty}.$ (61)
Let $f_{t}$ be the corresponding policy whose value function is $V_{A}^{f_{t}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}v_{t}$. Here, we assume that $\gamma_{A}>0$. For the case of $\gamma_{A}=0$, Statement (c) is an immediate consequence of Statement (d) because $v_{\infty}\in\mathcal{PV}$ and $\mathcal{PV}\subseteq\partial\mathcal{V}$.
Subsequently, we show $V_{A}^{f^{*}{\dagger}}\succcurlyeq Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})\succcurlyeq\bar{v}_{\infty}$ using Equation 60. In the proof of Statement (c) of Lemma 6.3, for all $s\in\mathcal{S}$,
@@
Therefore, we have $Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})\succcurlyeq\bar{v}$. Because $\bar{v}\succ v_{\infty}\succcurlyeq V_{A}^{f_{t}{\dagger}}$, it implies that $Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})\succcurlyeq V_{A}^{f_{t}{\dagger}}$. From Statement (a) of Lemma 6.3, we have $V_{A}^{f^{*}{\dagger}}\succcurlyeq Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})$.
@@
where we used Equation 61 for the last inequality. It implies $\mathcal{L}[v_{t+1}]>\mathcal{L}[v_{\infty}]$, which contradicts to the fact that $v_{\infty}\succcurlyeq v_{t+1}$. Therefore, $\min_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\right)$ for all $v\in\mathcal{V}$.
@@
Suppose that $v_{\infty}\in\mathcal{V}\setminus\partial\mathcal{V}$ (i.e., $v_{\infty}$ is an interior of $\mathcal{V}$). Then, there exists an $r>0$ such that =$\mathcal{E}=\{v\in\mathcal{F}_{\mathcal{S}}\mid\lVert v-v_{\infty}\rVert_{\infty}\leq r\}\subseteq\mathcal{V}$. Let $v_{\infty}^{r}$ be such that $v_{\infty}^{r}(s)=v_{\infty}(s)+r$ for all $s\in\mathcal{S}$. Then, $v_{\infty}^{r}\in\mathcal{E}$. Moreover, because $v_{\infty}^{r}\in\mathcal{V}$, there exists a policy $f_{\infty}^{r}$ such that $V_{A}^{f_{\infty}^{r}{\dagger}}=v_{\infty}^{r}$.
We lead to the contradiction to $\min_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\right)$. Using Equation 60, we obtain
@@
where the derivation of Equation 69a is due to the definition of $\mathcal{W}_{1-\epsilon}(f_{t})$ in Equation 9 and $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f{\dagger}}$. This implies $\mathcal{L}[v_{t+1}]>\mathcal{L}[v_{\infty}]$, which contradicts to $v_{\infty}\succcurlyeq v_{t+1}$. Therefore, $v_{\infty}\in\mathcal{PV}$. ∎
Appendix K Proposed Algorithm with Backtracking
@@
Algorithm 1 may not output a PO policy because the terminate condition in line 5 of Algorithm 1 is the necessary condition for PO policies. In this case, Lemma 6.5 can be used to determine whether the output policy is a PO policy: provided (i) or (ii) in Lemma 6.5 holds for all $f\in\mathcal{W}_{A}$ with the output policy $f_{t}$ and its state value function $V_{A}^{f_{t}{\dagger}}$, $f_{t}$ is a PO policy. Notably, this judgment is not perfect because the policy may be determined not to be a PO policy when it really is a PO policy.
@@
We proposed selecting another policy $f^{\prime}\in\mathcal{W}(f_{t-1})$ that has not been selected thus far and restarting the for loop with $f_{t}\leftarrow f^{\prime}$, which we call ”Backtracking”, to deal with the case where the output policy $f_{t}(t<M)$ is determined not to be a PO policy. If there is no policy left in $\mathcal{W}(f_{t-1})$, then it backtracks one by one, as $\mathcal{W}(f_{t-2}),\mathcal{W}(f_{t-3}),\dots$, until it reaches the set with policies that have never been selected. This method requires that $\mathcal{W}(f_{t})$ is saved at each iteration.
@@
The entire proposed algorithm including the PO-policy judgement and the backtracking is shown in Algorithm 2. $W$ is a list to save each $\mathcal{W}(f_{t})$ and $M$ is the maximum number of iterations. This algorithm is designed such that the number of times to compute $V_{A}^{f_{t}{\dagger}}$ in line 5 and $\mathcal{W}_{\succcurlyeq}(f_{t})$ in line 7 is at most $M$. If the PO-policy judgment in line 9 does not return True during $M$ iterations, the algorithm outputs the policy that maximizes the objective $\sum_{s\in\mathcal{S}}\alpha_{s}V_{A}^{f_{t}{\dagger}}(s)$ among the ones that satisfy the necessary condition for PO policies thus far and $f_{M}$ (see lines 12, 25, and 26).
Algorithm 2 PIPOP with Backtracking
@@
9:        if $f_{t}$ satisfies (i) or (ii) of Lemma 6.5 for all $f\in\mathcal{W}_{A}$ then
@@
Appendix L Relation between the method of Zhang et al. (2020a) and Bucarey et al. (2022)
Zhang et al.Zhang et al. (2020a) defined the optimal value functions as $V_{i}^{*}(i\in\{A,B\})$, which are the solution of the following equationsZhang et al. (2020a): let $a\in\mathcal{A},b\in\mathcal{B},\gamma\in[0,1)$, and for all $s\in\mathcal{S}$,
@@
$V_{i}^{*}(s)=Stackelberg_{i}(Q_{A}^{*}(s),Q_{B}^{*}(s)),$ (71)
where $Q_{i}^{*}(s)$ is the Q-value $Q_{i}(a,b,s)$ of an action pair $(a,b)\in\mathcal{A}\times\mathcal{B}$ in state $s$. Given the fixed $s$, $Q_{i}^{*}(b,a,s)$ can be viewed as the payoff of $(a,b)$ for the agent $i$; thus, $Q_{i}^{*}(s)$ can be viewed as the payoff table of agent $i$ in the state $s$. $Stackelberg_{i}(Q_{A}^{*}(s),Q_{B}^{*}(s))$ is the payoff for agent $i$ in the Stackelberg equilibrium of the normal-form game represented by the payoff tables $(Q_{A}^{*}(s),Q_{B}^{*}(s))$. Therefore, Equation 71 can be written as
@@
By substituting Equation 70 to Equation 72, we obtain
@@
Meanwhile, the operator used in Bucarey et al.Bucarey et al. (2022) is given by
$(Tv)_{i}(s)$ $=r_{i}(s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{b}))}\left[v_{i}(s^{\prime})\right],$ (74a)
$\mathrm{where~{}}$ $R_{A}(s,v):=\operatornamewithlimits{argmax}_{f_{s}\in\Delta(\mathcal{A})}r_{A}(s,f_{s},R_{B}(s,f_{s},v_{B}))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},R_{B}(s,f_{s},v_{B}))}\left[v_{A}(s^{\prime})\right],$ (74b)
$R_{B}(s,f_{s},v_{B}):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f_{s},b)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},b)}\left[v_{B}(s^{\prime})\right].$ (74c)
Notably, $\gamma=\gamma_{A}=\gamma_{B}$. Let us consider only the deterministic leader policies in the fixed point of Equation 74 and replace $\operatornamewithlimits{argmax}_{f_{s}\in\Delta(\mathcal{A})}$ to $\operatornamewithlimits{argmax}_{a\in\mathcal{A}}$ in the definition of $R_{A}$. Then, letting $(V_{A},V_{B})$ be the fixed point of $T$, it holds from Equation 74, for all $s\in\mathcal{S}$, which is expressed as
$V_{i}(s)$ $=r_{i}(s,R_{A}(s,V),R_{B}(s,R_{A}(s,V),V_{B}))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,V),R_{B}(s,R_{A}(s,V),V_{B}))}\left[V_{i}(s^{\prime})\right],$ (75a)
$\mathrm{where~{}}$ $R_{A}(s,V):=\operatornamewithlimits{argmax}_{a\in\mathcal{A}}r_{A}(s,a,R_{B}(s,a,V_{B}))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a,R_{B}(s,a,V_{B}))}\left[V_{A}(s^{\prime})\right],$ (75b)
$R_{B}(s,a,V_{B}):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,a,b)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a,b)}\left[V_{B}(s^{\prime})\right].$ (75c)
Comparing these with Equation 73, the definition of $(V_{A}^{*},V_{B}^{*})$ by Equation 73 is equivalent to the definition of $(V_{A},V_{B})$ by Equation 75. It follows that the fixed point in Zhang et al.Zhang et al. (2020a) is the same in Bucarey et al.Bucarey et al. (2022) when the leader policies in the fixed point are restricted to deterministic stationary policies.
By applying the discussion in the proof of Theorem 5.1 to $(V_{A},V_{B})$ defined by Equation 75, we obtain
@@
for all $s\in\mathcal{S}$. Then, the problem of the method of Bucarey et al. discussed in Section 5.1 also occurs to $(V_{A},V_{B})$: $V_{A}$ does not coincide with the state value function of the deterministic SSE policies since $R_{B}(f,V_{B})$ is not guaranteed to be the best response against any $f\in\mathcal{W}_{A}^{d}$.
Appendix M Application: Policy Teaching by Intervention to the Transitions
In single-agent MDPs, the policy aimed by the agent can be changed by making changes to the agent’s observations, reward signals, and transition probabilities, among others. For example, in an MDP $(\mathcal{S},\mathcal{A},p,\rho,r,\gamma)$, the changes in the observations and the transition probabilities are represented by replacement to another transition function $p^{\prime}$, which changes the agent’s aim from the optimal policies of $(\mathcal{S},\mathcal{A},p,\rho,r,\gamma)$ to those of $(\mathcal{S},\mathcal{A},p^{\prime},\rho,r,\gamma)$. This is the theoretical foundation of poisoning attacks against RL Behzadan & Munir (2017); Huang & Zhu (2019); Zhang et al. (2020b); Rakhsha et al. (2020); Sun et al. (2020) and policy teaching Zhang & Parkes (2008); Zhang et al. (2009), which aims to guide the agent’s policy learning.
SSGs can be viewed as a model of intervention in transition probabilities by the leader where the follower is the victim of the attack or the teaching. Even if the follower takes the same action at each state, the transition can be shifted by the leader changing its actions because the transition probability depends on the actions of all the agents. Therefore, suppose the follower cannot observe both the attendance and the actions of the leader, then the follower’s learning can be represented as learning the optimal policies of the single-agent MDP $(\mathcal{S},\mathcal{A}_{B},p^{f},\rho,\hat{r}_{B},\gamma_{B})$, where $f\in\mathcal{W}_{A}$ is the leader’s policy, $p^{f}(s^{\prime}|s,b):=\mathbb{E}_{a\sim f(s)}[\hat{p}(s^{\prime}|s,a,b)]$, and $\hat{r}_{B}:\mathcal{S}\times\mathcal{B}\to\mathbb{R}$. Notably, $R_{B}^{*}(f)$ is its optimal policy. Then, the problem of guiding the follower’s aim to the leader’s objective is expressed as
@@
where $\hat{r}_{A}$ is the leader’s reward function that represents the desired follower’s behavior intended by the leader. $f^{*}$ is the optimal solution of Equation 77, and the state-action sequences generated by the corresponding follower’s optimal policy $R_{B}^{*}(f)$ maximize not only the follower’s expected cumulative discounted reward but also that of the leader’s. This can be viewed as the follower attaining the leader’s desired behavior when the follower achieves its own learning goal .
The optimal solution of Equation 77 is given by the PO policy $f$ that maximizes $\mathcal{L}[V_{A}^{f{\dagger}}]=\sum_{s\in\mathcal{S}}\rho(s)V_{A}^{f{\dagger}}(s)$. Therefore, our proposed approach can be applied to this problem when the transition function is known and the follower’s best response can be computed. Although the optimality of the obtained solution is not guaranteed unless $\gamma_{A}=0$, our analysis guarantees the monotone improvement of the leader’s policy, which is of great importance in practice. In contrast, existing methods trying to obtain the SSE policy may be inadequate because it is usual that the follower is not myopic and hence the SSE policy does not necessarily exist.

### v2

@@ 7 Limitation and Future Work
$\mathcal{L}[v_{\infty}]-\mathcal{L}[v_{t}]$ $<(1-\epsilon)(\mathcal{L}[\bar{v}]-\mathcal{L}[v_{\infty}]),$ (44)
where $\bar{v}(s):=v_{\infty}(s)+(1-\delta)(\tau-\gamma_{A})\lVert v^{*}-v_{\infty}\rVert_{\infty}.$ Let $f_{t}$ be the corresponding policy whose value function is $V_{A}^{f_{t}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}v_{t}$. Here, we assume that $\gamma_{A}>0$. For the case of $\gamma_{A}=0$, Statement (c) is an immediate consequence of Statement (d) because $v_{\infty}\in\mathcal{PV}$ and $\mathcal{PV}\subseteq\partial\mathcal{V}$.
Subsequently, we show $V_{A}^{f^{*}{\dagger}}\succcurlyeq Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})\succcurlyeq\bar{v}_{\infty}$ using Equation˜43. In the proof of Statement (c) of Theorem˜6.3, for all $s\in\mathcal{S}$,
@@
Therefore, we have $Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})\succcurlyeq\bar{v}$. Because $\bar{v}\succ v_{\infty}\succcurlyeq V_{A}^{f_{t}{\dagger}}$, it implies that $Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})\succcurlyeq V_{A}^{f_{t}{\dagger}}$. From Statement (a) of Theorem˜6.3, we have $V_{A}^{f^{*}{\dagger}}\succcurlyeq Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})$.
@@
where we used Equation˜44 for the last inequality. It implies $\mathcal{L}[v_{t+1}]>\mathcal{L}[v_{\infty}]$, which contradicts to the fact that $v_{\infty}\succcurlyeq v_{t+1}$. Therefore, we have $\min_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\right)$ for all $v\in\mathcal{V}$.
@@
Suppose that $v_{\infty}\in\mathcal{V}\setminus\partial\mathcal{V}$ (i.e., $v_{\infty}$ is an interior of $\mathcal{V}$). Then, there exists an $r>0$ such that $\mathcal{E}=\{v\in\mathcal{F}_{\mathcal{S}}\mid\lVert v-v_{\infty}\rVert_{\infty}\leq r\}\subseteq\mathcal{V}$. Let $v_{\infty}^{r}$ be such that $v_{\infty}^{r}(s)=v_{\infty}(s)+r$ for all $s\in\mathcal{S}$. Then, $v_{\infty}^{r}\in\mathcal{E}$. Moreover, because $v_{\infty}^{r}\in\mathcal{V}$, there exists a policy $f_{\infty}^{r}$ such that $V_{A}^{f_{\infty}^{r}{\dagger}}=v_{\infty}^{r}$.
We lead to the contradiction to $\min_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\right)$. Using Equation˜43, we obtain
@@
where the derivation of Equation˜52a is due to the definition of $\mathcal{W}_{1-\epsilon}(f_{t})$ in Equation˜7 and $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f{\dagger}}$. This implies $\mathcal{L}[v_{t+1}]>\mathcal{L}[v_{\infty}]$, which contradicts to $v_{\infty}\succcurlyeq v_{t+1}$. Therefore, $v_{\infty}\in\mathcal{PV}$.
Appendix G Splitting the Search Space
The proposed algorithm in Algorithm˜1 can be impractical when the leader’s policy space $\mathcal{W}_{A}$ is prohibitively large to search for the next policy in Equation˜6 and Equation˜8. To avoid the exhaustive search, we propose a splitting strategy for the search space to obtain improved policies from each subspace efficiently.
We first split $\mathcal{W}_{A}$ into disjoint sets $\mathcal{W}_{1}\cup\dots\cup\mathcal{W}_{K}$, where the best response is $R_{B}^{*}(f)=g_{i}^{*}\in\mathcal{W}_{B}^{d}$ for $f\in\mathcal{W}_{i}$. Because the set of deterministic policies, $\mathcal{W}_{B}^{d}$, is finite, we can always find such a separation. Additionally, we argue that the number of deterministic policies that form the best response is rather limited, enabling us to enumerate them easily.
We consider replacing (8) with locating a Pareto optimal $f$ with respect to $Q_{A}^{f_{t}\dagger}(\cdot,f)$. Because (8) also locates Pareto optimal $f$, this replacement corresponds to selecting a different $\mathcal{L}$ at each $t$.
This is realized in two steps. First, we locate a policy satisfying $Q_{A}^{f_{t}\dagger}(\cdot,f)\succcurlyeq V_{A}^{f_{t}\dagger}$. It is nontrivial as the best response $R_{B}^{*}$ changes as $f$ changes. However, if we limit our attention to a subset $\mathcal{W}_{k}$ for some $k$, the best response does not change with $f\in\mathcal{W}_{k}$. Then, locating $f\in\mathcal{W}_{k}$ satisfying $Q_{A}^{f_{t}\dagger}(\cdot,f)\succcurlyeq V_{A}^{f_{t}\dagger}$ is casted as the following constrained optimization problem:
$\min_{z,f\in\mathcal{W}_{k}}z\quad\text{s.t.}\quad V_{A}^{f_{t}{\dagger}}(s)-Q_{A}^{f_{t}\dagger}(s,f)\leqslant z\quad\text{for all }s\in\mathcal{S},$ (53)
where $Q_{A}^{f_{t}\dagger}(s,f)$ is differentiable with respect to $f$ and is easily computable as the best response is constant. Therefore, it will be solved by, e.g., a gradient-based solver with projection onto $\mathcal{W}_{k}$. A solution with $z\leq 0$ satisfies $Q_{A}^{f_{t}\dagger}(\cdot,f)\succcurlyeq V_{A}^{f_{t}\dagger}$. Such a solution must exist in $\mathcal{W}_{k}$ for some $k$ because $f_{t}$ satisfies this condition.
Once we find such a solution, denoted as $(z_{t},f_{t}^{\prime})$, we find a Pareto optimal policy for $Q_{A}^{f_{t}\dagger}(\cdot,f)$ in $\mathcal{W}_{k}\cap\mathcal{W}_{\succcurlyeq}(f_{t})$ by improving $f_{t}^{\prime}$. It is done by performing a Pareto ascent method with projection,
$f\leftarrow\Pi_{\mathcal{W}_{k}}(f+\eta\Delta),\quad\text{where}\quad\Delta^{\top}\nabla Q_{A}^{f_{t}\dagger}(s,f)\geqslant 0\quad\text{for all }s\in\mathcal{S},$ (54)
or its variants, e.g., Harada et al. (2006), with $f_{t}$ as a feasible initial solution of this step. If we find $f_{t}^{*}$ such that $Q_{A}^{f_{t}\dagger}(\cdot,f)\succ V_{A}^{f_{t}\dagger}$, we let $f_{t+1}=f_{t}^{*}$. These steps will be continued until a predefined termination criterion is satisfied, or no policy strictly dominating the current policy is found. The whole algorithm is shown in Algorithm˜2.
Algorithm 2 Practical Pareto-Optimal Policy Iteration
0: Maximum number of iterations $M$.
1: Split the leader’s policy space $\mathcal{W}_{A}$ into disjoint sets $\mathcal{W}_{1}\cup\dots\cup\mathcal{W}_{K}$, where the best response is $R_{B}^{*}(f)=g_{i}^{*}\in\mathcal{W}_{B}^{d}$ for $f\in\mathcal{W}_{i}$
2: Randomly initialize $f_{0}\in\mathcal{W}_{A}$ and compute $g_{0}=R_{B}^{*}(f_{0})$
3:for $t=0$ to M1M-1 do
4:  Compute $V_{A}^{f_{t}{\dagger}}$ by repeatedly applying $T_{A}^{f_{t}R_{B}^{*}(f_{t})}$
5:  Let $\mathcal{I}_{t}=\{1,\dots,K\}$
6:  while $\lvert\mathcal{I}_{t}\rvert>0$ do
7:   Randomly select $k\in\mathcal{I}_{t}$
8:   Solve (53) to obtain $z_{t},f_{t}^{\prime}$
9:   if $z_{t}\leq 0$ then
10:    Perform the Pareto ascent (54) to obtain $f_{t}^{*}$
11:    Let $f_{t+1}=f_{t}^{*}$ and break if $Q_{A}^{f_{t}\dagger}(\cdot,f_{t}^{*})\succ V_{A}^{f_{t}{\dagger}}$
12:   end if
13:   Update $\mathcal{I}_{t}\leftarrow\mathcal{I}_{t}\setminus\{k\}$
14:  end while
15:   Let $f_{M}=\dots=f_{t+1}=f_{t}$ and break if $\mathcal{I}_{t}=\emptyset$
16:end for
16: A stationary policy $f_{M}$
Appendix H Backtracking
@@
Algorithm˜1 may not output a PO policy because the terminate condition in line 5 of Algorithm˜1 is the necessary condition for PO policies. In this case, Theorem˜6.5 can be used to determine whether the output policy is a PO policy: provided (i) or (ii) in Theorem˜6.5 holds for all $f\in\mathcal{W}_{A}$ with the output policy $f_{t}$ and its state value function $V_{A}^{f_{t}{\dagger}}$, $f_{t}$ is a PO policy. Notably, this judgment is not perfect because the policy may be determined not to be a PO policy when it really is a PO policy.
@@
We proposed selecting another policy $f^{\prime}\in\mathcal{W}(f_{t-1})$ that has not been selected thus far and restarting the for loop with $f_{t}\leftarrow f^{\prime}$, which we call "Backtracking", to deal with the case where the output policy $f_{t}(t<M)$ is determined not to be a PO policy. If there is no policy left in $\mathcal{W}(f_{t-1})$, then it backtracks one by one, as $\mathcal{W}(f_{t-2}),\mathcal{W}(f_{t-3}),\dots$, until it reaches the set with policies that have never been selected. This method requires that $\mathcal{W}(f_{t})$ is saved at each iteration.
@@
The entire proposed algorithm with the PO-policy judgement and the backtracking is shown in Algorithm˜3. $W$ is a list to save each $\mathcal{W}(f_{t})$ and $M$ is the maximum number of iterations. This algorithm is designed such that the number of times to compute $V_{A}^{f_{t}{\dagger}}$ in line 5 and $\mathcal{W}_{\succcurlyeq}(f_{t})$ in line 7 is at most $M$. If the PO-policy judgment in line 9 does not return True during $M$ iterations, the algorithm outputs the policy that maximizes the objective $\sum_{s\in\mathcal{S}}\alpha_{s}V_{A}^{f_{t}{\dagger}}(s)$ among the ones that satisfy the necessary condition for PO policies thus far and $f_{M}$ (see lines 12, 25, and 26).
Algorithm 3 Pareto-Optimal Policy Iteration with Backtracking
@@
9:   if $f_{t}$ satisfies (i) or (ii) of Theorem˜6.5 for all $f\in\mathcal{W}_{A}$ then
@@
Appendix I Relation between the methods of Zhang et al. (2020) and Bucarey et al. (2022)
Zhang et al. (2020a) defined the optimal value functions as $V_{i}^{*}(i\in\{A,B\})$, which are the solution of the following equations (Zhang et al., 2020a): let $a\in\mathcal{A},b\in\mathcal{B},\gamma\in[0,1)$, and for all $s\in\mathcal{S}$,
@@
$V_{i}^{*}(s)=\textsc{Stackelberg}_{i}(Q_{A}^{*}(s),Q_{B}^{*}(s)),$ (56)
where $Q_{i}^{*}(s)$ is the Q-value $Q_{i}(a,b,s)$ of an action pair $(a,b)\in\mathcal{A}\times\mathcal{B}$ in state $s$. Given the fixed $s$, $Q_{i}^{*}(b,a,s)$ can be viewed as the payoff of $(a,b)$ for the agent $i$; thus, $Q_{i}^{*}(s)$ can be viewed as the payoff table of agent $i$ in the state $s$. $\textsc{Stackelberg}_{i}(Q_{A}^{*}(s),Q_{B}^{*}(s))$ is the payoff for agent $i$ in the Stackelberg equilibrium of the normal-form game represented by the payoff tables $(Q_{A}^{*}(s),Q_{B}^{*}(s))$. Therefore, Equation˜56 can be written as
@@
By substituting Equation˜55 to Equation˜57, we obtain
@@
Meanwhile, the operator used in Bucarey et al. (2022) is given by
[…]
$R_{B}(s,f_{s},v_{B}):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f_{s},b)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},b)}\left[v_{B}(s^{\prime})\right].$ (59g)
Notably, $\gamma=\gamma_{A}=\gamma_{B}$. Let us consider only the deterministic leader policies in the fixed point of Equation˜59 and replace $\operatornamewithlimits{argmax}_{f_{s}\in\Delta(\mathcal{A})}$ with $\operatornamewithlimits{argmax}_{a\in\mathcal{A}}$ in the definition of $R_{A}$. Then, letting $(V_{A},V_{B})$ be the fixed point of $T$, it holds from Equation˜59, for all $s\in\mathcal{S}$, which is expressed as
[…]
Comparing these with Equation˜58, the definition of $(V_{A}^{*},V_{B}^{*})$ by Equation˜58 is equivalent to the definition of $(V_{A},V_{B})$ by Equation˜60. It follows that the fixed point in Zhang et al. (2020) is the same in Bucarey et al. (2022) when the leader policies in the fixed point are restricted to deterministic stationary policies.
By applying the discussion in the proof of Theorem˜5.1 to $(V_{A},V_{B})$ defined by Equation˜60, we obtain
@@
for all $s\in\mathcal{S}$. Then, the problem of the method of Bucarey et al. discussed in Section˜5 also occurs to $(V_{A},V_{B})$: $V_{A}$ does not coincide with the state value function of the deterministic SSE policies since $R_{B}(f,V_{B})$ is not guaranteed to be the best response against any $f\in\mathcal{W}_{A}^{d}$.
Appendix J Application: Policy Teaching by Intervention to the Transitions
In single-agent MDPs, the policy aimed by the agent can be changed by making changes to the agent’s observations, reward signals, and transition probabilities, among others. For example, in an MDP $(\mathcal{S},\mathcal{A},p,\rho,r,\gamma)$, the changes in the observations and the transition probabilities are represented by replacement to another transition function $p^{\prime}$, which changes the agent’s aim from the optimal policies of $(\mathcal{S},\mathcal{A},p,\rho,r,\gamma)$ to those of $(\mathcal{S},\mathcal{A},p^{\prime},\rho,r,\gamma)$. This is the theoretical foundation of poisoning attacks against RL (Behzadan and Munir, 2017; Huang and Zhu, 2019; Zhang et al., 2020b; Rakhsha et al., 2020; Sun et al., 2020) and policy teaching (Zhang and Parkes, 2008; Zhang et al., 2009), which aims to guide the agent’s policy learning.
SSGs can be viewed as a model of intervention in transition probabilities by the leader, where the follower is the victim of the attack or the teaching. Even if the follower takes the same action at each state, the transition can be shifted by the leader changing its actions because the transition probability depends on the actions of all the agents. Therefore, suppose the follower cannot observe both the attendance and the actions of the leader, then the follower’s learning can be represented as learning the optimal policies of the single-agent MDP $(\mathcal{S},\mathcal{A}_{B},p^{f},\rho,\hat{r}_{B},\gamma_{B})$, where $f\in\mathcal{W}_{A}$ is the leader’s policy, $p^{f}(s^{\prime}|s,b):=\mathbb{E}_{a\sim f(s)}[\hat{p}(s^{\prime}|s,a,b)]$, and $\hat{r}_{B}:\mathcal{S}\times\mathcal{B}\to\mathbb{R}$. Notably, $R_{B}^{*}(f)$ is its optimal policy. Then, the problem of guiding the follower’s aim to the leader’s objective is expressed as
@@
where $\hat{r}_{A}$ is the leader’s reward function that represents the desired follower’s behavior intended by the leader. $f^{*}$ is the optimal solution of Equation˜62, and the state-action sequences generated by the corresponding follower’s optimal policy $R_{B}^{*}(f)$ maximize not only the follower’s expected cumulative discounted reward but also that of the leader’s. This can be viewed as the follower attaining the leader’s desired behavior when the follower achieves its own learning goal.
The optimal solution of Equation˜62 is given by the PO policy $f$ that maximizes $\mathcal{L}[V_{A}^{f{\dagger}}]=\sum_{s\in\mathcal{S}}\rho(s)V_{A}^{f{\dagger}}(s)$. Therefore, our proposed approach can be applied to this problem when the transition function is known and the follower’s best response can be computed. Although the optimality of the obtained solution is not guaranteed unless $\gamma_{A}=0$, our analysis guarantees the monotone improvement of the leader’s policy, which is of great importance in practice. In contrast, existing methods trying to obtain the SSE policy may be inadequate because it is usual that the follower is not myopic, and hence, the SSE policy does not necessarily exist.